OpenStudy (anonymous):

What is the domain of the function? 3x^2-27 / x^2+7x+12 =y

OpenStudy (anonymous):

here take 3 outside that's 3(x^2-9) that's 3(x^2-3^2) = 3(x+3)(x-3) in denominator x^2+7x+12 =(x+3)(x+4) after solving u wil get 3(x-3)/x+4) = y now find the domain domain = value in which function is defined

OpenStudy (anonymous):

I did something wrong because i had 3(x-9) / x+4-y=0

OpenStudy (anonymous):

haha you have to simplify equation and then find domain it gets easy so whats your domain

OpenStudy (anonymous):

I am extremely confused is it 0?

OpenStudy (anonymous):

is it -3 and -4